9+1=? vgvhsfbjdjhjsfhkghjksd...........................................................................................................................................................................................................................................................................................

9+1=?
vgvhsfbjdjhjsfhkghjksd...........................................................................................................................................................................................................................................................................................

Grade:10

7 Answers

aziz
6 years ago
See the answer to this question is very simple.
9+1=10
Divyanshu Kumar
6 years ago
Answer is 10.we going to bed now so I`m not going to be with you and I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to the power of the African American Idol Finale Countdown I have to go to
sahil
6 years ago
it is very very simple the answer is 10 as 9+1=10 i think you wanted to ask any other thing if yes type again
Sujit Kumar
6 years ago
Your question is 9+1=?
The answer is 10 unless I misunderstood your question or you were trying to ask something else. Please do not post irrelevant questions on askIITians. This question might be taken down soon. 
 
praveen kumar
6 years ago
are you testing the quality of answers in askiitians. it’s very  simple 9+1=10
thank you. 
please ask the doubts that you have we are ready to solve that.
thanks again.
Amit
6 years ago
9+1=10
Vivek
6 years ago
9+1=10 acc to mathematics its very simple.its just addition. I think yiu have some other difficulty which you should clearly ask it.